# FlowPath: Learning Data-Driven Manifolds with Invertible Flows for Robust Irregularly-sampled Time Series Classification

### v1

@@ Abstract
Extended versionhttps://arxiv.org/abs/
@@ Introduction
Many real-world systems evolve continuously over time. In contrast, observations of these systems are typically sparse, irregular, and incomplete (choi_doctor_2016; hossain_method_2020; oh_stable_2024). Such data are referred to as Irregularly-sampled time series (ISTS), characterized by non-uniform sampling intervals and missing observations over time. While discrete-time models such as GRUs (chung_empirical_2014) have been adapted to handle irregularity using decay mechanisms (che_recurrent_2018), they remain fundamentally discrete approximations. Neural Differential Equations (NDEs) address this limitation by modeling time as a continuous variable, with Neural Ordinary Differential Equations (Neural ODEs) as a foundational case (chen_neural_2018; lu_beyond_2018).
For ISTS, Neural Controlled Differential Equations (Neural CDEs) extend this idea by modeling system dynamics as driven by a continuous control path derived from the input (kidger_neural_2020). Unlike Neural ODEs that evolve from an initial state (rubanova_latent_2019), Neural CDEs respond directly to the full input stream. However, the control path is typically constructed via fixed interpolation methods, which impose predefined geometric assumptions. This choice has been shown to affect performance substantially and remains a key limitation of the Neural CDE framework (morrill_choice_2022).
In this paper, we propose FlowPath, a framework that models the control path using an invertible neural flow (dinh_density_2017; papamakarios_normalizing_2021). This replaces the fixed interpolant with a learned, data-adaptive transformation that preserves invertibility. Such flows serve as universal approximators of diffeomorphisms (teshima_coupling-based_2020; ishikawa_universal_2023), allowing the learned path to be smooth and information-preserving. We formalize the FlowPath architecture, analyze its theoretical properties, and evaluate it on 18 benchmark datasets and real-world datasets. Our results demonstrate that learning a structurally-constrained, data-driven manifold leads to statistically significant improvements in classification accuracy over baselines using either fixed interpolants or non-invertible learned paths.
@@
Traditional Recurrent Neural Networks (RNNs) assume fixed step sizes (goodfellow_deep_2016), and adaptations such as GRU-D (che_recurrent_2018) incorporate missingness indicators or time-lag decay to handle irregular sampling, but remain discrete-time approximations that can still struggle with large sampling gaps or nonuniform intervals. Thus, a conventional RNN or GRU-D processes data at sampled time points $\{t_{k}\}$. An RNN state $\bm{h}_{k}\in\mathbb{R}^{d_{h}}$ typically evolves via:
@@
To overcome the limitations of purely discrete-time methods, chen_neural_2018 introduced Neural ODEs, casting hidden state evolution as a continuous function of time,
@@
The state $\bm{z}(t)\in\mathbb{R}^{d_{z}}$ in a Neural ODE is obtained by integrating an initial value $\bm{z}(0)$ from 0 to $T$. While this continuous viewpoint elegantly handles irregular time points, its strict reliance on the initial value problem makes it difficult to incorporate new observations arriving at later times without restarting or extending the integration (rubanova_latent_2019; kidger_neural_2022).
Neural CDEs (kidger_neural_2020) overcome this limitation by representing the input time series as a control path $X(t)\in\mathbb{R}^{d_{x}}$, thereby enabling the hidden state to evolve continuously in response to incoming observations. The hidden state dynamics are governed by a Riemann–Stieltjes integral with respect to the control path:
@@
Control path $X(t)$ is typically constructed from observed data using fixed interpolation methods such as linear or cubic splines (morrill_neural_2021). This introduces a structural prior that may not align with the true dynamics. Empirical results show that interpolation choice is a sensitive hyperparameter with significant impact on performance, and no single method is optimal across tasks (morrill_choice_2022).
Recent studies replace the fixed interpolant with a learned path from unconstrained networks, such as encoder-decoders (jhin_exit_2022; jhin_learnable_2023). Others apply attention mechanisms within the Neural CDE framework to reweight temporal contributions (jhin_attentive_2024). Further work focuses on improving the expressivity of the dynamics by signature transformation (morrill_neural_2021). These approaches enhance flexibility but may introduce instability when the learned path lacks structural constraints.
@@
bilos_neural_2021 propose modeling ODE solutions with an invertible transformation, circumventing numerical solvers by directly parameterizing a diffeomorphism in time. Such invertible mappings have been extensively studied in normalizing flows (papamakarios_normalizing_2021), where tractable Jacobians enable density estimation or generative sampling. Formally, Neural Flow (bilos_neural_2021) is designed to learn an invertible mapping $F(t,\bm{x}(0);\theta_{F})$ that directly parameterizes the solution of an ODE or dynamical process. Given an implicit dynamical system governed by,
@@
Crucially, $F(\cdot)$ is designed to be a diffeomorphism in $\bm{x}$, ensuring that mapping local volumes is always invertible. This property allows Neural Flow to avoid the need for standard numerical ODE solvers and to guarantee the model’s Jacobian is tractable. While originally aimed at continuous normalizing flows for generative tasks, the concept of applying invertible flows to dynamic time series modeling has grown (bilos_neural_2021; oh_dualdynamics_2025).
@@
Let $F\colon[0,T]\times\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{x}}$ be a neural network parameterized by $\theta_{F}$ such that the mapping $F$ is a diffeomorphism in its second argument (bilos_neural_2021). We explicitly define the learnable control path $\Phi:[0,T]\to\mathbb{R}^{d_{x}}$ as:
@@
We define a hidden state $\bm{z}(t)\in\mathbb{R}^{d_{z}}$ governed by the Riemann–Stieltjes integral as suggested by kidger_neural_2020:
$\bm{z}(t)=\bm{z}(0)+\int_{0}^{t}f\bigl(\tau,\bm{z}(\tau);\theta_{f}\bigr)\,\mathrm{d}\Phi(t).$ (2)
@@
with $\bm{z}(0)=h\bigl(\bm{x}(0)\bigr)$, and $h\colon\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{z}}$ an embedding layer. In Eq. (3), the dot notation $\dot{\bm{z}}(t)$ and $\dot{\Phi}(t)$ denote the time derivatives, i.e., $\frac{d\bm{z}(t)}{dt}$ and $\frac{d\Phi(t)}{dt}$, respectively. The smoothness assumptions on $\Phi$ guarantee Lipschitz continuity in time, while invertibility enforces distribution-preserving transformations (papamakarios_normalizing_2021; bilos_neural_2021). Since $\Phi$ is a diffeomorphism, the learned path $t\;\mapsto\;\Phi(t)$ admits well-defined inverses and Jacobians, which strengthen theoretical guarantees such as existence, uniqueness, and universal approximation for continuous-time processes. In practice, numerical solvers treat $\dot{\Phi}(t)$ as a known function (output of a neural net) and integrate $\dot{\bm{z}}(t)$ forward in time, following the ODE form in Eq. (3).
@@
We parameterize the control path using an invertible neural network based on normalizing flows, which provide a flexible framework for learning smooth, bijective transformations (papamakarios_normalizing_2021). Recent theoretical results show that such architectures can approximate any diffeomorphism under suitable conditions (teshima_coupling-based_2020; ishikawa_universal_2023). This ensures that the learned path $\Phi(t)$ remains invertible and smooth by design, reducing the risk of instability often seen in unconstrained path parameterizations.
Within the FlowPath framework, $\Phi(t)=F(t,\bm{x}(0);\theta_{F})$ acts as a trainable control path that modulates the evolution of the latent state $\bm{z}(t)$. The function $F$ is designed as an invertible neural network, drawing inspiration from normalizing flows (papamakarios_normalizing_2021) and specifically neural flows that parameterize diffeomorphisms (bilos_neural_2021). This design ensures that $\Phi(t)$ is a smooth transformation of its inputs, providing FlowPath with enhanced flexibility to model complex temporal dependencies.
@@
Under the FlowPath framework in Eq. (2), let $\Phi(t)$ be a $C^{1}$-diffeomorphism, and $f$ be Lipschitz continuous in $\bm{z}$}. Then the probability density $p(z(t))$ of the latent state evolves according to
@@
This property, while crucial for continuous normalizing flows in generative modeling (papamakarios_normalizing_2021), also benefits our classification task by preserving class‐separating structures in latent space, thereby improving robustness to irregular sampling and missing data.
@@
The universal approximation properties of Neural ODEs and Neural Flows have been extensively studied and established in prior work (lin_resnet_2018; li_deep_2022; teshima_coupling-based_2020; ishikawa_universal_2023). Building upon these results, one can check that FlowPath is a universal approximator under appropriate regularity conditions. Under the assumption that $\Phi$ is a $C^{1}$-diffeomorphism and $f$ is Lipschitz, Neural ODEs are universal approximators for any smooth target trajectory $u$ in the uniform norm. Thus, one can choose $f$ such that within any desired precision $\epsilon^{\prime}$
@@
This result, derived via Rademacher complexity arguments (bartlett_spectrally-normalized_2017; chen_generalization_2020), indicates that, as the number of training samples $n$ grows, the empirical loss of a FlowPath model closely approximates its true expected loss on unseen data.
@@
We design experiments to evaluate both the classification accuracy and robustness of our model under varying degrees of missing observations. Specifically, we consider 18 diverse time series classification datasets from UEA & UCR Data Repository, referenced in (bagnall_uea_2018; h_a_dau_ucr_2019), utilizing the sktime Python library (loning_sktime_2019). We followed the experimental setup and dataset selection described by oh_dualdynamics_2025. The original paper considered three categories: ‘Motion & HAR’, ‘Electrocardiogram (ECG) & Electroencephalogram (EEG)’, and ‘Sensor’ domains, as summarized in the supplementary material.
@@
We utilized the conventional RNN (medsker_recurrent_1999), LSTM (s_hochreiter_long_1997), and GRU (chung_empirical_2014), with mean imputation. Furthermore, we employed several modifications of the GRU, specifically GRU-$\Delta t$ (choi_doctor_2016), GRU-D (che_recurrent_2018), which have been tailored to efficiently handle ISTS. Additionally, we considered models that are based on the principles of Neural ODEs, such as GRU-ODE (brouwer_gru-ode-bayes_2019), ODE-RNN (rubanova_latent_2019) and ODE-LSTM (lechner_learning_2020). Moreover, recent advancements of Neural CDEs were also included in the evaluation, such as Neural CDE (kidger_neural_2020), Neural RDE (morrill_neural_2021), ANCDE (jhin_attentive_2024), EXIT (jhin_exit_2022), LEAP (jhin_learnable_2023), and DualDynamics (oh_dualdynamics_2025). Lastly, we included naïve Neural Flow (bilos_neural_2021) with different configurations: ResNet, GRU, and Coupling Flow.
@@
(d) Flow-based invertible path
@@
To evaluate robustness on a real-world task, we used the PAMAP2 dataset (a_reiss_introducing_2012), a challenging benchmark for sensor-based Human Activity Recognition (HAR). Each participant was equipped with three sensors positioned on the wrist, chest, and ankle, as well as a heart rate monitor. The dataset comprises 5333 segments spanning 8 daily living activities, captured using 17 distinct sensor modalities.
We followed the exact preprocessing protocol, data splits, and sensor dropout methodology proposed by zhang_graph-guided_2022. To simulate irregularity, we randomly removed 60% of data points from each segment, using the same random mask across all experiments. After that, we generated scenarios by randomly removing a subset of input sensors with dropout rates ranging from 10% to 50%. This approach increases the challenge of the problem, as the omitted sensors are randomly selected for each sample. For every test sample, a subset of sensors was designated as missing, and all observations from these sensors were replaced with zeros.
@@
We compare the proposed FlowPath with the following state-of-the-art baselines, including Transformer (vaswani_attention_2017), Trans-mean (Transformer with mean imputation), GRU-D (chen_neural_2018), SeFT (horn_set_2020), mTAND (shukla_multi-time_2021), Raindrop (zhang_graph-guided_2022), and Neural CDE (kidger_neural_2020).
For the original dataset without sensor dropout, we have included additional benchmarks such as DGM2-O (wu_dynamic_2021), IP-Net (shukla_interpolation-prediction_2019), MTGNN (wu_connecting_2020), TITD (ji_titd_2025), and CoFormer (y_wei_compatible_2023). For fair comparison, we used performance results as reported in the original literature.
@@
Settings Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
Regular 95.0 $\pm$0.4 93.9 $\pm$0.4 94.6 $\pm$0.5 95.6 $\pm$0.2 94.8 $\pm$0.2
@@
Table 5: Classification performance on the PAMAP2 dataset with all sensors under irregular sampling.
@@
Settings Methods Accuracy Precision Recall F1 Score
10%
Missing
Transformer 60.9 $\pm$12.8 58.4 $\pm$18.4 59.1 $\pm$16.2 56.9 $\pm$18.9
@@
We further evaluated model performance on a medical classification benchmark on the PhysioNet Sepsis dataset (reyna_early_2020), which includes 40,335 patients and 34 temporal variables. Following kidger_neural_2020, we compared models with and without observation intensity (OI), an index reflecting patient severity, and measured classification performance by AUROC.
Benchmark models and reported performance from kidger_neural_2020, oh_dualdynamics_2025, and oh_comprehensive_2025. We also include variants of stable Neural SDEs (oh_stable_2024), including Neural Langevin-type SDE (LSDE), Neural Linear Noise SDE (LNSDE), and Neural Geometric SDE (GSDE). For further implementation details and benchmark methods, please refer to the original studies.
Table 7: AUROC Comparison on PhysioNet Sepsis

### v2

@@ Abstract
Extended versionhttps://arxiv.org/abs/2511.10841
@@ Introduction
Many real-world systems evolve continuously over time. In contrast, observations of these systems are typically sparse, irregular, and incomplete (Choi et al. 2016; Hossain et al. 2020; Oh et al. 2024). Such data are referred to as irregularly-sampled time series (ISTS), characterized by non-uniform sampling intervals and missing observations over time. While discrete-time models such as GRUs (Chung et al. 2014) have been adapted to handle irregularity using decay mechanisms (Che et al. 2018), they remain fundamentally discrete approximations. Neural Differential Equations (NDEs) address this limitation by modeling time as a continuous variable, with Neural Ordinary Differential Equations (Neural ODEs) as a foundational case (Chen et al. 2018; Lu et al. 2018).
For ISTS, Neural Controlled Differential Equations (Neural CDEs) extend this idea by modeling system dynamics as driven by a continuous control path derived from the input (Kidger et al. 2020). Unlike Neural ODEs that evolve from an initial state (Rubanova et al. 2019), Neural CDEs respond directly to the full input stream. However, the control path is typically constructed via fixed interpolation methods, which impose predefined geometric assumptions. This choice has been shown to affect performance substantially and remains a key limitation of the Neural CDE framework (Morrill et al. 2022).
In this paper, we propose FlowPath, a framework that models the control path using an invertible neural flow (Dinh et al. 2017; Papamakarios et al. 2021). This replaces the fixed interpolant with a learned, data-adaptive transformation that preserves invertibility. Such flows serve as universal approximators of diffeomorphisms (Teshima et al. 2020; Ishikawa et al. 2023), allowing the learned path to be smooth and information-preserving. We formalize the FlowPath architecture, analyze its theoretical properties, and evaluate it on 18 benchmark datasets and real-world datasets. Our results demonstrate that learning a structurally-constrained, data-driven manifold leads to statistically significant improvements in classification accuracy over baselines using either fixed interpolants or non-invertible learned paths.
@@
Traditional Recurrent Neural Networks (RNNs) assume fixed step sizes (Goodfellow et al. 2016), and adaptations such as GRU-D (Che et al. 2018) incorporate missingness indicators or time-lag decay to handle irregular sampling, but remain discrete-time approximations that can still struggle with large sampling gaps or nonuniform intervals. Thus, a conventional RNN or GRU-D processes data at sampled time points $\{t_{k}\}$. An RNN state $\bm{h}_{k}\in\mathbb{R}^{d_{h}}$ typically evolves via:
@@
To overcome the limitations of purely discrete-time methods, Chen et al. (2018) introduced Neural ODEs, casting hidden state evolution as a continuous function of time,
@@
The state $\bm{z}(t)\in\mathbb{R}^{d_{z}}$ in a Neural ODE is obtained by integrating an initial value $\bm{z}(0)$ from 0 to $T$. While this continuous viewpoint elegantly handles irregular time points, its strict reliance on the initial value problem makes it difficult to incorporate new observations arriving at later times without restarting or extending the integration (Rubanova et al. 2019; Kidger 2022).
Neural CDEs (Kidger et al. 2020) overcome this limitation by representing the input time series as a control path $X(t)\in\mathbb{R}^{d_{x}}$, thereby enabling the hidden state to evolve continuously in response to incoming observations. The hidden state dynamics are governed by a Riemann–Stieltjes integral with respect to the control path:
@@
Control path $X(t)$ is typically constructed from observed data using fixed interpolation methods such as linear or cubic splines (Morrill et al. 2021). This introduces a structural prior that may not align with the true dynamics. Empirical results show that interpolation choice is a sensitive hyperparameter with significant impact on performance, and no single method is optimal across tasks (Morrill et al. 2022).
Recent studies replace the fixed interpolant with a learned path from unconstrained networks, such as encoder-decoders (Jhin et al. 2022, 2023). Others apply attention mechanisms within the Neural CDE framework to reweight temporal contributions (Jhin et al. 2024). Further work focuses on improving the expressivity of the dynamics by signature transformation (Morrill et al. 2021). These approaches enhance flexibility but may introduce instability when the learned path lacks structural constraints.
@@
Bilos et al. (2021) propose modeling ODE solutions with an invertible transformation, circumventing numerical solvers by directly parameterizing a diffeomorphism in time. Such invertible mappings have been extensively studied in normalizing flows (Papamakarios et al. 2021), where tractable Jacobians enable density estimation or generative sampling. Formally, Neural Flow (Bilos et al. 2021) is designed to learn an invertible mapping $F(t,\bm{x}(0);\theta_{F})$ that directly parameterizes the solution of an ODE or dynamical process. Given an implicit dynamical system governed by,
@@
Crucially, $F(\cdot)$ is designed to be a diffeomorphism in $\bm{x}$, ensuring that mapping local volumes is always invertible. This property allows Neural Flow to avoid the need for standard numerical ODE solvers and to guarantee the model’s Jacobian is tractable. While originally aimed at continuous normalizing flows for generative tasks, the concept of applying invertible flows to dynamic time series modeling has grown (Bilos et al. 2021; Oh et al. 2025b).
@@
Let $F\colon[0,T]\times\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{x}}$ be a neural network parameterized by $\theta_{F}$ such that the mapping $F$ is a diffeomorphism in its second argument (Bilos et al. 2021). We explicitly define the learnable control path $\Phi:[0,T]\to\mathbb{R}^{d_{x}}$ as:
@@
We define a hidden state $\bm{z}(t)\in\mathbb{R}^{d_{z}}$ governed by the Riemann–Stieltjes integral as suggested by Kidger et al. (2020):
$\bm{z}(t)=\bm{z}(0)+\int_{0}^{t}f\bigl(\tau,\bm{z}(\tau);\theta_{f}\bigr)\,\mathrm{d}\Phi(\tau).$ (2)
@@
with $\bm{z}(0)=h\bigl(\bm{x}(0)\bigr)$, and $h\colon\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{z}}$ an embedding layer. In Eq. (3), the dot notation $\dot{\bm{z}}(t)$ and $\dot{\Phi}(t)$ denote the time derivatives, i.e., $\frac{d\bm{z}(t)}{dt}$ and $\frac{d\Phi(t)}{dt}$, respectively. The smoothness assumptions on $\Phi$ guarantee Lipschitz continuity in time, while invertibility enforces distribution-preserving transformations (Papamakarios et al. 2021; Bilos et al. 2021). Since $\Phi$ is a diffeomorphism, the learned path $t\;\mapsto\;\Phi(t)$ admits well-defined inverses and Jacobians, which strengthen theoretical guarantees such as existence, uniqueness, and universal approximation for continuous-time processes. In practice, numerical solvers treat $\dot{\Phi}(t)$ as a known function (output of a neural net) and integrate $\dot{\bm{z}}(t)$ forward in time, following the ODE form in Eq. (3).
@@
We parameterize the control path using an invertible neural network based on normalizing flows, which provide a flexible framework for learning smooth, bijective transformations (Papamakarios et al. 2021). Recent theoretical results show that such architectures can approximate any diffeomorphism under suitable conditions (Teshima et al. 2020; Ishikawa et al. 2023). This ensures that the learned path $\Phi(t)$ remains invertible and smooth by design, reducing the risk of instability often seen in unconstrained path parameterizations.
Within the FlowPath framework, $\Phi(t)=F(t,\bm{x}(0);\theta_{F})$ acts as a trainable control path that modulates the evolution of the latent state $\bm{z}(t)$. The function $F$ is designed as an invertible neural network, drawing inspiration from normalizing flows (Papamakarios et al. 2021) and specifically neural flows that parameterize diffeomorphisms (Bilos et al. 2021). This design ensures that $\Phi(t)$ is a smooth transformation of its inputs, providing FlowPath with enhanced flexibility to model complex temporal dependencies.
@@
Under the FlowPath framework in Eq. (2), let $\Phi(t)$ be a $C^{1}$-diffeomorphism, and $f$ be Lipschitz continuous in $\bm{z}$}. Then the probability density $p({\bm{z}}(t))$ of the latent state evolves according to
@@
This property, while crucial for continuous normalizing flows in generative modeling (Papamakarios et al. 2021), also benefits our classification task by preserving class‐separating structures in latent space, thereby improving robustness to irregular sampling and missing data.
@@
The universal approximation properties of Neural ODEs and Neural Flows have been extensively studied and established in prior work (Lin and Jegelka 2018; Li et al. 2022; Teshima et al. 2020; Ishikawa et al. 2023). Building upon these results, one can check that FlowPath is a universal approximator under appropriate regularity conditions. Under the assumption that $\Phi$ is a $C^{1}$-diffeomorphism and $f$ is Lipschitz, Neural ODEs are universal approximators for any smooth target trajectory $u$ in the uniform norm. Thus, one can choose $f$ such that within any desired precision $\epsilon^{\prime}$
@@
This result, derived via Rademacher complexity arguments (Bartlett et al. 2017; Chen et al. 2020), indicates that, as the number of training samples $n$ grows, the empirical loss of a FlowPath model closely approximates its true expected loss on unseen data.
@@
We design experiments to evaluate both the classification accuracy and robustness of our model under varying degrees of missing observations. Specifically, we consider 18 diverse time series classification datasets from UEA & UCR Data Repository, referenced in (Bagnall et al. 2018; H. A. Dau et al. 2019), utilizing the sktime Python library (Löning et al. 2019). We followed the experimental setup and dataset selection described by Oh et al. (2025b). The original paper considered three categories: ‘Motion & Human Activity Recognition (HAR)’, ‘Electrocardiogram (ECG) & Electroencephalogram (EEG)’, and ‘Sensor’ domains, as summarized in the supplementary material.
@@
We utilized the conventional RNN (Medsker and Jain 1999), LSTM (S. Hochreiter and J. Schmidhuber 1997), and GRU (Chung et al. 2014), with mean imputation. Furthermore, we employed several modifications of the GRU, specifically GRU-$\Delta t$ (Choi et al. 2016), GRU-D (Che et al. 2018), which have been tailored to efficiently handle ISTS. Additionally, we considered models that are based on the principles of Neural ODEs, such as GRU-ODE (Brouwer et al. 2019), ODE-RNN (Rubanova et al. 2019) and ODE-LSTM (Lechner and Hasani 2020). Moreover, recent advancements of Neural CDEs were also included in the evaluation, such as Neural CDE (Kidger et al. 2020), Neural RDE (Morrill et al. 2021), ANCDE (Jhin et al. 2024), EXIT (Jhin et al. 2022), LEAP (Jhin et al. 2023), and DualDynamics (Oh et al. 2025b). Lastly, we included naïve Neural Flow (Bilos et al. 2021) with different configurations: ResNet, GRU, and Coupling Flow.
@@
(d) Flow-based
invertible path
@@
To evaluate robustness on a real-world task, we used the PAMAP2 dataset (A. Reiss and D. Stricker 2012), a challenging benchmark for sensor-based Human Activity Recognition (HAR). Each participant was equipped with three sensors positioned on the wrist, chest, and ankle, as well as a heart rate monitor. The dataset comprises 5333 segments spanning 8 daily living activities, captured using 17 distinct sensor modalities.
We followed the exact preprocessing protocol, data splits, and sensor dropout methodology proposed by Zhang et al. (2022). To simulate irregularity, we randomly removed 60% of data points from each segment, using the same random mask across all experiments. After that, we generated scenarios by randomly removing a subset of input sensors with dropout rates ranging from 10% to 50%. This approach increases the challenge of the problem, as the omitted sensors are randomly selected for each sample. For every test sample, a subset of sensors was designated as missing, and all observations from these sensors were replaced with zeros.
@@
We compare the proposed FlowPath with the following state-of-the-art baselines, including Transformer (Vaswani et al. 2017), Trans-mean (Transformer with mean imputation), GRU-D (Chen et al. 2018), SeFT (Horn et al. 2020), mTAND (Shukla and Marlin 2021), Raindrop (Zhang et al. 2022), and Neural CDE (Kidger et al. 2020).
For the original dataset without sensor dropout, we have included additional benchmarks such as DGM2-O (Wu et al. 2021), IP-Net (Shukla and Marlin 2019), MTGNN (Wu et al. 2020), TITD (Ji et al. 2025), and CoFormer (Y. Wei et al. 2023). For fair comparison, we used performance results as reported in the original literature.
@@
Sensor Dropout Neural CDE FlowPath
Spline MLP ResNet GRU Coupling
0% 95.0 $\pm$0.4 93.9 $\pm$0.4 94.6 $\pm$0.5 95.6 $\pm$0.2 94.8 $\pm$0.2
10% 87.3 $\pm$1.3 86.4 $\pm$1.0 87.2 $\pm$0.3 88.3 $\pm$0.4 87.4 $\pm$0.3
20% 76.8 $\pm$0.6 75.1 $\pm$0.8 79.3 $\pm$1.0 79.1 $\pm$1.9 77.8 $\pm$0.6
30% 67.1 $\pm$1.9 68.2 $\pm$0.8 71.4 $\pm$0.7 70.0 $\pm$0.7 70.2 $\pm$1.0
40% 57.7 $\pm$1.6 58.8 $\pm$2.8 65.9 $\pm$1.9 59.6 $\pm$1.0 60.8 $\pm$1.1
50% 51.6 $\pm$1.0 50.7 $\pm$2.0 57.9 $\pm$0.2 54.1 $\pm$1.0 54.2 $\pm$3.1
@@
Table 5: Classification performance on the PAMAP2 dataset with all sensors under original irregular sampling.
@@
Methods Accuracy Precision Recall F1 Score
10% Missing Transformer 60.9 $\pm$12.8 58.4 $\pm$18.4 59.1 $\pm$16.2 56.9 $\pm$18.9
@@
We further evaluated model performance on a medical classification benchmark on the PhysioNet Sepsis dataset (Reyna et al. 2020), which includes 40,335 patients and 34 temporal variables. Following Kidger et al. (2020), we compared models with and without observation intensity (OI), an index reflecting patient severity, and measured classification performance by AUROC.
Benchmark models and reported performance from Kidger et al. (2020), Oh et al. (2025b), and Oh et al. (2025a). We also include variants of stable Neural SDEs (Oh et al. 2024), including Neural Langevin-type SDE (LSDE), Neural Linear Noise SDE (LNSDE), and Neural Geometric SDE (GSDE). For implementation details and benchmark methods, please consult the original studies.